OpenStudy (anonymous):

Determine all solutions of the differential equation y`= -9y Use the variables C and x as appropriate.

OpenStudy (bahrom7893):

dy/dx = -9y dy/y = -9dx ln|y| = -9x + C e^ln|y| = e^(-9x+C) y = e^C*e^(-9x) y = Ae^(-9x)

OpenStudy (bahrom7893):

as e to some constant is still a constant then u just write that as A

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

yeh np

OpenStudy (anonymous):

haha, click on become a fan, really want to get to the next level, haha, I like it. You deserve it bahrom, nice work. Wish I had someone to help me out like that the first time I saw diffeq ; )